One fair coin is tossed 15 times, what is the probability of getting 3 heads out of 15 tossing experiments?

Answer by stanbon(75887)

One fair coin is tossed 15 times, what is the probability of getting 3 heads out of 15 tossing experiments?
-------
Binomial Problem with n = 15 ; p(heads) = 1/2 = p(tails)
---------------
P(x = 3 heads) = 15C3(1/2)^3*(1/2)^12
----
= 455/2^15
---
= 455/32768
---
= 0.0139
==============
Cheers,
Stan H.
===============